### v1

# Ting En Lam1,2, Yuhan Chen1,3, Elston Tan1,4, Eric Peh1, Ruirui Chen1, Paritosh Parmar1, Basura Fernando1

Abstract
Causal video question answering (QA) has garnered increasing interest, yet existing datasets often lack depth in causal reasoning analysis. To address this gap, we capitalize on the unique properties of cartoons and construct CausalChaos!, a novel, challenging causal Why-QA dataset built upon the iconic “Tom and Jerry" cartoon series. With thoughtful questions and multi-level answers, our dataset contains much longer causal chains embedded in dynamic interactions and visuals, at the same time principles of animation allows animators to create well-defined, unambiguous causal relationships. These factors allow models to solve more challenging, yet well-defined causal relationships. We also introduce hard negative mining, including CausalConfusion version. While models perform well, there is much room for improvement, especially, on open-ended answers. We identify more advanced/explicit causal relationship modeling & joint modeling of vision and language as the immediate areas for future efforts to focus upon. Along with the other complementary datasets, our new challenging dataset will pave the way for these developments in the field. We will release our dataset, codes, and models to help future efforts in this domain.
Figure 1: (Left) Examples of causal questions about characters’ actions from our CausalChaos! dataset—based on Tom & Jerry cartoon series. Q: question; A: answer; E: explanation. Please view in Adobe Reader to play the embedded videos for better explanation. (Right, top) Illustration of causal chain, scene changes. Linking multiple clues/cues embedded in different scenes to solve causal relationships pose a challenge for VideoQA models. (Right, bottom) Being a synthetic medium allows animators to leverage Principles of Animation to stylize the visuals & motions to disentangle/highlight key content of the scene to create well-defined/unambiguous and effectively communicated cause-and-effect relationships. The interplay of these factors allow models to focus on solving complex, yet, well-defined, unambiguous causal relationships.
1 Introduction
Causal video question answering has been gaining increasing interest [52, 28, 47, 17]. However, our analysis, as presented in Figs.1, 2, and in § 2, indicates that existing causal QA datasets may often require surface-level understanding or involve more of a simple word substitution in the QA pairs, rather than causal reasoning (e.g., “Q. Why are the cars on the street not moving? A. Parked."). Consequently, these datasets do not demand true causal reasoning analysis. This gap leaves a dearth of reliable benchmarks for evaluating the causal reasoning capabilities of video understanding models. Hence, in this study, we concentrate on developing a rigorous and challenging causal Video-QA dataset. We aim to create a high-quality resource that can effectively test and facilitate the advancement of sophisticated causal models.
Drawing on the established benefits of cartoons with cognitive processes among children [38, 39, 30, 53, 40], we leverage the renowned series [Uncaptioned image] to create a novel and demanding causal Why-QA dataset called CausalChaos!. The key characteristics of CausalChaos! are as follows:
  • As illustrated in Figs.1, 2, our emphasis lies in crafting thoughtful questions paired with multi-layered answers, aiming to cultivate a deeper and more comprehensive grasp of causal relationships. Compared to existing datasets, CausalChaos! presents extended chains of causality [^1] A causal chain is a sequence of events/actions in which each step influences or leads to the next, creating a cause-and-effect relationship. It illustrates how one event/action produces a subsequent event/outcome., as highlighted in Fig.2 and discussed in § 2.
  • Video clips in our dataset feature frequent scene and shot changes [^2] A scene change or shot change refers to the transition from one visual setting or perspective to another. It typically involves shifting focus to a new location, action, or character within the storyline, often signaled by techniques such as cuts & fades.. The links or intermediary causes within the causal chains are often dispersed across various scenes. Consequently, models are challenged to exert greater cognitive effort in connecting multiple events (scenes) and identifying intermediate causes to comprehend the “why" questions related to the clips.
  • Despite the complexity and length of the causal chains, they are distinctly delineated, devoid of ambiguity, and communicated effectively using principles of animation [42] such as staging and exaggeration. This deliberate design allows models to concentrate on deciphering causal relationships.
  • CausalChaos! introduces an added layer of complexity by necessitating the modeling of actions at various levels of granularity—ranging from sweeping, large-scale movements to nuanced, subtle actions, such as interpreting emotional cues through facial expressions.
  • Our dataset demands a diverse range of reasoning skills, encompassing deductive, spatial, emotional reasoning, and more, as outlined in Fig.2 and discussed in §3.
We evaluated various state-of-the-art Video-QA models including recent multimodal instruction tuning models and found that our dataset remains the most challenging causal QA dataset. Evaluating on our dataset, helped us reveal that models generally jumped to conclusions based on partial evidence, instead of taking into consideration full set of evidence; models failed to engage in true causal reasoning, trying to select answers based on shortcuts like object/action-noun/verb matching. Based on this, we identified more advanced/explicit causal relationship modeling and joint modeling vision and language as the immediate areas for future efforts to focus upon. Further, we show that similar to how cartoons help children better connect cause and effects, they can help Video-QA models as well. We incorporated our dataset with the NextQA [47], a real-world dataset and found that it brings some improvements in why questions. What is more, incorporating our dataset brings improvement on non-Why questions as well.
Overall, we believe that our CausalChaos! dataset contributes challenges that span the entire Video-QA pipeline, from deciphering intricate videos to processing complex questions and discerning between nuanced answers, stimulating research in many areas like video processing, causal reasoning, language modeling and joint modeling. Along with the other complementary datasets, such as [47, 17, 19, 21, 46], our new challenging dataset will pave the way for these developments in the field. We advocate for leveraging our dataset sourced from cartoons to inform the model design process. Our dataset better reflects the challenges Video-QA models may face in the real world, such as longer causal chains and frequent scene transitions in visual streams. Nevertheless, we caution against blindly transferring or deploying entire models/weights trained on our dataset to real-world applications.
Dataset
Multilevel
Answers
Causal
Chain Len
Scene
Reasoning
Spectrum
Multigranular
Actions
NextQA [47] 1 1 narrow limited
CausalVidQA [17] 1 1 narrow limited
IntentQA [19] 1 1 narrow limited
Ours 2.7 4 wide ubiquitous
[…]
[loop,autoplay,poster=1,width=]3Figs/vvn_frs_pdf/000015
Figure 2: (a) Types of reasoning demanded by our CausalChaos! dataset. Reasoning types: DR-deductive reasoning; IR-inductive; SR-spatial; CR-causal; CT-critical thinking; ER-emotional; AR-abductive; TR-temporal; None-no reasoning required as per the human subjects. None is undesirable, and tend to indicate that questions are less challenging. (b) Comparison among CausalChaos! and existing causal videoQA datasets. (c) Qualitative comparison between CausalChaos! and NextQA datasets. CausalChaos! Answers and Explanations give detailed information regarding the actual cause-and-effect relationships, motivations, emotions covering wide range of reasoning types. Note that, here, with the help of human annotators, we have temporally cropped videos to retain only the relevant parts from NextQA dataset videos; otherwise, raw videos are longer resulting in unintended problem of temporal localization for models.
2 Related Work
VideoQA Datasets and Benchmarks. To drive the progress in VideoQA, researchers have developed various datasets with distinct focuses and contributions—see the survey [59]. In the following, we mainly compare and contrast our work from prior art, with a focus on causal QA literature. We provide a detailed discussion on Video QA datasets in the Supplementary material.
Early VideoQA datasets
such as MovieFIB [26], MSVD-QA [48], MSRVTT-QA [48], YouTube2Text [10], open-ended QA [58], Zeng et al. [56], and Video Context-QA [61] relied on video captions or descriptions to automatically generate questions and corresponding answers, primarily focusing on object and action recognition. However, they lacked depth in understanding complex interactions and causal relationships within videos. TGIF-QA [13, 14] and ActivityNet-QA [54], while covering short and long web videos respectively, still fell short in capturing comprehensive reasoning, particularly in object interactions and causal relationships. Social-IQ [55] addressed human social behaviour in videos through multimodal cues but lacked broader context and visual reasoning capabilities. AGQA [9] focused on spatio-temporal scene understanding, but did not cover explanatory "why" questions. STAR [46], a situation reasoning dataset, extended to prediction and feasibility questions but also lacked explanatory inquiries.
Causal QA datasets.
Our work is inspired from a few datasets: CLEVRER [52, 28], NextQA [47], CausalVidQA [17], IntentQA [19] that cover causal reasoning. However, they have several limitations that our dataset aims to fill:
  • Some are limited by their scope, e.g., [52, 28] only consider collision events of simple inanimate objects such as cuboids; as a result, lack: actors, their characteristics such as emotions, intentions, lack diverse language descriptions.
  • Some datasets (e.g., [47, 17]) do not contain any precise temporal annotations, which inadvertently involves temporal localization & may create a false sense of difficulty, while the questions may not be challenging. We alleviate this problem by providing precise temporal annotations and focusing on designing truely complex and challenging QA.
  • Some datasets lack complexity in questions [47, 17, 19]. Based on human studies, we found that a majority of causal QA in these datasets were flagged as not requiring any notable reasoning (Figure 2). We also attempted to quantify the complexity in reasoning by computing the lengths of causal chains involved in their QAs vs. ours using GPT-4 [32]. Causal chains in [47, 17, 19] have an average length of 1. In comparison, our dataset has much longer causal chains—average length 2.7—posing a challenge for models in connecting multiple events or cues. Our dataset also demands wide spectrum of types of reasoning (2(a)).
  • Scenes in [52, 28, 47, 17, 19] are less dynamic. For example, they mostly involve a single scene. In comparison, our dataset, on average, has about 4 scene changes. Rapid scene changes & dynamic interactions pose a challenge to VideoQA models when modeling cause-&-effect relationships over longer causal chains by linking context and cues embedded in different scenes (more details in subsection 3.2). While related challenges have been acknowledged in other computer vision problems [34], they remain unexplored in the realm of VideoQA.
  • Current datasets do not emphasize on hard incorrect option selection. As such, models may not require to do causal reasoning, but rather they can exploit shortcuts like object/action-noun/verb matching in vision-language spaces to correctly answer. Thus, we develop hard negative selection strategies.
  • These datasets contain a single-level answer to ‘Why’-questions. On the other hand, we richly annotate our dataset to provide two-level answers to ‘Why’-questions. Quality of answers is further enhanced, and are more informative as a consequence of the more complex questions.
VideoQA Tasks/Models.
VideoQA generally uses two protocols: 1) multichoice QA (MCQA) & 2) open-ended answer generation (OEAG). MCQA involves choosing the correct option from the multiple provided options. MCQA models [2, 8, 6, 16, 15, 7] fuse Video, question, & answer representations & process them using a multi-layer perceptron, which ultimately predicts the correct answer. On the other hand, in much more challenging OEAG, VideoQA models generate answers in natural language [2, 51, 6, 15]. This is typically accomplished by decoding the fused video & question representations into a sequence of words (answer). Recently, vision-language and multimodal instruction tuning models [7, 18, 57, 21] have shown strong performance on video understanding tasks. We evaluate & report the performance of all these models on our dataset. Further discussion on individual models can be found in Supplementary material.
3 CausalChaos! Dataset
In this section, we detail the construction of our challenging VideoQA dataset focused on causal reasoning. We provide insights into the video source, annotation process, and rigorous quality checks employed to ensure detailed annotations of high quality. Further, we discuss the unique attributes of our dataset that make it a valuable resource for the causal action/VideoQA task.
3.1 Dataset Construction
Video source.
To exclusively focus on visual information for reasoning, we selected the timeless silent cartoon series, ‘Tom & Jerry’ (1940), spanning over 6 seasons and encompassing a total of 161 episodes. Silent cartoons that do not contain modalities other than vision (such as dialogues) can foster the development of visual reasoning capabilities in children. Similarly, we hypothesize that using silent cartoons for VideoQA can allow for a focus on visual reasoning. The Tom & Jerry series is also an ideal source for constructing a causal reasoning dataset, as it provides abundant segments showcasing diverse cause-&-effect relationships. A concurrent work [22] on causal image generation has also leveraged Tom & Jerry cartoon.
Annotations.
We collect the following annotations.
$\bullet$ Questions. The questions are crafted to capture the why or reasoning behind the actions of characters in Tom & Jerry cartoons. To cultivate comprehensive and deeper video understanding capabilities in models, we formulated thoughtful causal questions where cause & effect are connected by longer causal chains. Questions also extend beyond explicit visual cues, encompassing gestures & expressions, to delve into the characters’ underlying intentions and goals.
The annotators would generate questions while watching the video for the first time to ensure that they did not have prior knowledge of the scene, which mimics how models would assess unseen clips. Subsequently, they would rewatch the videos multiple times to generate more critical thinking questions.
While generating questions, annotators watch the videos without audio to ensure that the dataset does not contain audio cues to focus on visual reasoning.
$\bullet$ Multi-level Answers. The first-level or the Primary answer, represents a shallower, literal or direct cause or form of response. It is accompanied by a deeper form of Explanation, which considers the broader context of the scene, the thoughts, feelings, intentions of characters, & their actions. This deeper explanation also takes into account potential consequences & provides further reasoning to support the primary answer. It includes reasons & additional details to comprehensively address the question. Examples shown in Figure 1 & Supplementary material.
$\bullet$ Temporal annotations. For each QAE set the Start &End frame numbers of the video are recorded such that the entire scene including frames that offer context & details supporting the reasons behind the actions is covered.
In summary, each dataset sample is annotated with: {start frame, end frame, Question, Answer, Explanation}.
Quality Check.
To ensure the quality and reliability of the dataset, we implemented a two-stage quality check process. In the first stage, quality checkers conducted thorough assessments of episodes they did not annotate. Subsequently, in the second stage, quality checkers reviewed episodes for which they did not perform the first-stage quality check.
Multiple quality checkers were involved in the process, carefully reviewing and verifying curated question and answer sets for logical fallacies, timestamp inconsistencies, grammatical errors, and spelling mistakes. The process also involved checking and excluding annotations containing audio or text aspects that the model cannot comprehend. To prevent direct overwriting, any edits were flagged in a different font color for further discussion.
The inconsistencies or errors that were flagged out were resolved through discussions and consensus among the checkers and respective annotators. In these discussions, the team referred to the episode synopses available on Tom and Jerry Wiki [43], an online Tom and Jerry encyclopedia. This external source served as a third-party opinion and facilitated fact-checking. By incorporating multiple perspectives, the discussions helped to provide a deeper level of analysis to generate a more challenging and comprehensive dataset.
Dataset Statistics.
Annotators meticulously watched all 161 cartoon episodes spanning 6 seasons, generating a total of 4945 richly annotated Question-Answer-Explanation sets for the dataset. The inclusion of a substantial number of Question-Answer-Explanation sets across all seasons, covering a wide range of scenes, characters, and events, underscores the comprehensive nature of our dataset. This diversity establishes a robust foundation for understanding and answering questions related to the cartoon series. Word clouds corresponding to answers and explanations from our dataset are presented in Supplementary material.
3.2 Unique Characteristics of our dataset
In the following, we delve into the untapped unique properties of our dataset, which both complement existing video understanding datasets and present novel challenges to VideoQA, particularly from a causal reasoning perspective.
  1. 1.
    Multi-level answers. VideoQA datasets generally provide single-level answers. Single-level answers might contain less information. Multi-level answers are valuable because they offer a richer and more nuanced perspective, which leads to deeper insights—considering not only the immediate circumstances but also underlying psychological, social, and personal factors that contributed to characters’ decisions. By employing multi-level answers alongside thoughtful questions, we can cultivate deeper reasoning and analysis of characters’ actions, enabling the training and evaluation of models on complex question-answer pairs that demand a comprehensive understanding of longer causal chains. Further elaborated discussion on the importance of multi-level answers to ‘Why’-questions included in Supplementary material.
  2. 2.
    Grounded in diverse Visual & Motion information. The Tom and Jerry cartoon series is visually rich and dynamic with intricate scenes, actions, and character interactions varying across all episodes. We asked annotators to craft questions where the answer is grounded in the video. By grounding questions(-answers) in the visual input, the model is required to analyze the video to gain access to a broader range of details and cues/clues from the videos to provide meaningful answers. For example, the same question—Why was Jerry laughing?—appears in two different seasons (episodes S03E17, S05E13). However, the Explanations are different: 1) for S03E17, the Explanation is “Jerry found it funny that the seal stole the fish from Tom."; 2) while for S05E13, the Explanation is “Jerry was laughing at Tom who fell down the stairs and into the fountain." This concretely shows that VideoQA models need to understand visual information to correctly answer the questions. Grounding QAE set in vision is underscored by the presence of unusual situations in cartoons—typically not found in real-world cases. For example, Jerry removing Tom’s nose and fitting strawberry in place of Tom’s nose or Lion hushing Jerry. Due to this, models cannot answer just based on the correlations/biases in its training data, rather models ought to process and understand the contents thoroughly to answer correctly. This is crucial even in with large language models (LLMs)/embeddings, which are also susceptible to such frequency-based biases/hallucinations [20, 24, 60]. For example, Lion is closely related to killing in embeddings from LLMs; in a qualitative analysis, we found that models may tend to choose an answer containing ‘killing’ Jerry as the motive for Lion-character’s action, while in actuality the Lion-character was trying to quiet Jerry. Revealing such biases has been considered in other computer vision problems [4, 12], but is yet to be introduced in VideoQA. Our dataset offers a valuable opportunity to address this challenge. Additionally, understanding Tom & Jerry cartoons requires modeling of details and actions at varying levels of granularity—from sweeping movements to subtle emotional cues via facial expressions.
  3. 3.
    Focus on causal reasoning in visually dynamic scenes. Our dataset focuses on causal reasoning in Tom and Jerry cartoons, which have an abundance of dynamic and rapidly changing scenes, making it challenging for traditional video understanding models to track and comprehend the context effectively. To effectively model cause-and-effect relationships and answer causal questions, video understanding models must comprehend the temporal and contextual flow of events, linking them together amid a multitude of rapid scene changes and dynamic interactions. Furthermore, characters and objects may appear, disappear, and reappear, potentially limiting the model’s access to only partial observations. We term this as Dynamic Scene Linking (DSL). DSL is different from temporal modeling in traditional sense, which is typically concerned with modeling temporal evolution of events within a scene (more details in Supplementary material). Related challenges have been recognized in other computer vision problems [34, 45], but not yet in VideoQA. We hypothesize that humans effectively understand these cartoons by forming a mental ‘world model’ of the scene. This mental model aids in bridging gaps between discontinuous scenes and provides a coherent representation of the overall context. Incorporating such capabilities into VideoQA has the potential to enhance model performance. Notably, current VideoQA datasets often lack coverage of such complex scenarios, unlike our dataset.
  4. 4.
    More challenging and requiring cognitive effort. Understanding causal chains can help explain why things happen and how different factors influence each other. Identifying and solving longer causal chains in the process of answering causal (why) questions involve linking together multiple cues/clues, which requires significant cognitive efforts on the part of VideoQA models, especially, when the clues/cues are embedded in different scenes (described above) like in our dataset. Annotators were explicitly instructed to craft QAE pairs that require linking together multiple cues under multiple scene changes. It is difficult to quantify the complexity of causal questions. To the best of our knowledge, there are no existing approaches to accomplish this. To this end, we propose using Causal Chain Length as a way to measure the complexity of causal questions. Computing lengths of Causal Chains. We leverage GPT-4 [32] to compute the lengths of causal chains of 100 randomly chosen QAE pairs from our dataset and existing causal(-Why) VideoQA datasets [47, 17, 19]. Results are presented & compared in Figure 2. We observed that the QAE pairs in our dataset have significantly longer causal chains than existing QA datasets. This finding sheds light on the unique challenges posed by our dataset and underscores the importance of developing robust VideoQA models capable of handling complex causal reasoning tasks. Examples of causal chains and additional details on computing causal chain lengths provided in Supplementary material.
  5. 5.
    Leveraging principles of animation [42] (shown in Figure 1) such as timing, squash and stretch, anticipation, staging, and exaggeration aid in: 1) “highlighting" key movements, emotions, & storytelling; 2) Consequently, greatly aid in establishing clear cause-and-effect relationships and effectively communicating them. These principles can be thought of as spatiotemporal counterparts of caricatures, which exaggerate and manipulate facial and bodily features, and these have been shown to improve facial recognition rates [37, 11, 29]. In other domains of computer vision, some works [41, 50, 49, 5] acknowledge the advantages and devise methods to disentangle the area of interest, enabling more targeted processing and comprehension of image contents. Similarly, we hypothesize that the principles of animation can provide “hints/guidance", but the models still need to be able to leverage this guidance to solve causal relationships. Overall, causal relationships in our dataset are complex, longer, but at the same time they are unambiguous/well-defined using principles of animation, allowing models to focus on deciphering causal relationships.
[…]
    [loop,autoplay,poster=1,width=0.95]3Figs/tor_frs/fr_0015
    Figure 3: Examples of various types of reasoning required by our dataset. Please zoom in and view in AdobeReader to watch the embedded videos for better explanation.
  6. 6.
    Wide spectrum of types of reasoning required. Our dataset demands various types of reasoning like 1) Deductive; 2) Inductive; 3) Spatial; 4) Causal; 5) Critical thinking; 6) Emotion; 7) Abductive; and 8) Temporal. We have visualized examples demonstrating these reasoning types in Figure 3. It is also possible that a question may involve more than one type of reasoning. We conducted human studies to determine the types of reasoning required by our dataset and NextQA and observed that our dataset demands a wider range of reasoning than NextQA (results in 2(a)). More information and definitions of types of reasoning and human studies are provided in Supplementary material.
  7. 7.
    Character names used instead of general nouns. In our dataset, we employ character names rather than generic nouns, enhancing contextual understanding and knowledge representation. The utilization of character names necessitates character/actor resolution, encouraging Video QA models to better comprehend question context and character dynamics. This contextual understanding results in more precise and contextually fitting answers. This approach becomes particularly crucial in scenes with multiple instances of the same type of animals, such as several cats or mice. Relying on character names ensures that models are required to distinguish between individual characters and accurately identify who is interacting with whom. Failure to discern these interactions may lead to inaccurate responses.
3.3 Benchmarks
Question-types/Tasks.
Our dataset comprises of following types of questions.
  1. 1.
    Multiple Choice Question Answering. VideoQA models are provided with the question, the associated video footage, the correct answer, and four incorrect answers; the task for the models is to pick the correct answer.
  2. 2.
    Open-ended Answer Generation. The VideoQA models are provided only with the question and the video footage, and the models have to generate the answers in natural language.
Incorrect answer/option mining.
In designing MCQA, the answer options should be sufficiently distinct from both the correct answer & each other. However, they should also possess enough semantic similarities to challenge reasoning capabilities beyond common sense. This way models cannot answer correctly by taking shortcuts like simple object/action matching in vision-language spaces. Naive Negative mining or using randomly chosen answers from other questions as incorrect options is unproductive; it renders the correct answer too obvious, so we avoid this strategy. Rather, we introduce the following strategies.
  1. 1.
    Vanilla Hard Negative mining. A pre-trained Sentence-Bert was adopted in the measure of semantic similarity. For each question of interest, we first sample top-10 other questions based on the cosine-similarity measure of the Sentence-Bert embedding. From these 10 question candidates, the associated answers are then sampled, and the cosine-similarity measure is once again applied between the correct answer and the answer candidate. Using a similarity measure in the equation: $\alpha cossim(Q_{i},Q_{n})\cdot\beta cossim(A_{i},A_{n})$, where $\alpha$ and $\beta$ can be varied to adjust the weights coming from the questions or answers, the top 4 answers are included as the option candidates. To prevent candidates from being too similar or rephrasing the form of the answer, the measure heavily penalizes answers with a high similarity $cossim(A_{i},A_{n})$ by setting $\beta$ = 0.001 to reduce the overall score of the candidate.
  2. 2.
    Causal-Confusion Negative mining. We leverage the power of large language models to generate confusing incorrect options which have the opposite effect (e.g. Tom was hitting Jerry $\rightarrow$ Tom was not hitting Jerry) or the causal agents are swapped (e.g. Tom was hitting Jerry $\rightarrow$ Jerry was hitting Tom). These causally-opposite incorrect options play an important role in causal action question-answering datasets like ours. We hypothesize that the models might not be doing sufficient joint modeling of language and vision. As such, Video QA models might be able to answer based on matching actors and objects in questions and answers. However, when tested on Causally-Opposite incorrect options, simple question-answer matching might not be enough, and consequently, Video QA models may fail to answer correctly. We hypothesize that simple objects and actors might be modeled by VideoQA models, but relationships might not be as they require more dynamic features and as such are harder to model. We have provided examples from Causal-Confusion vs. Vanilla Hard Negative in the Supplementary material.
Design of train-val-test splits.
In general, we divide the dataset into a train-val-test split ratio of 70:15:15. We also use the following train/test strategies.
  1. 1.
    Seen-episode testing.
    1. (a)
      Uniformly Distributed Seen-episode testing (UD). In this case, dataset samples are randomly divided into train-val-test splits. No constraints such as the train samples should be drawn only from certain episodes or only from certain parts of certain episodes are imposed. Due to this, it is likely that uniformly distributed chunks/portions of all episodes from all seasons are seen during training. This way the VideoQA models can potentially have sparse storylines with them.
    2. (b)
      Consecutive Partially Seen-episode testing (PS). In this case, train set is constrained to be drawn from starting 70% video of each episode; val set is constrained to be drawn from the middle 70-85% video of each episode; while the test set is constrained to be drawn from the remaining last 15% of the video of each episode. This creates a scenario where the model has partially “seen" episodes during training stage and then does “educated guessing" when answering during testing stage.
  2. 2.
    Unseen-episode testing (UN). In this case, a certain 70% of the episodes are reserved for train set; 15% of the remaining episodes are reserved for val set; and the remaining 15% of the episodes are used to prepare the test set. So, testing is done on entirely novel, unseen episodes and storylines.
4 Experiments
We benchmark the performance of state-of-the-art VideoQA methods on our newly introduced causal action QA dataset on two tasks: 1) MCQA; 2) OEAG.
UD PS UN
Model A A+E A A+E A A+E
Chance 20.00 04.00 20.00 04.00 20.00 04.00
BlindQA [2] 29.38 13.07 26.51 11.54 25.13 11.02
EVQA [2] 29.38 13.48 31.32 13.32 27.82 14.65
CoMem [8] 32.08 13.88 26.10 09.89 23.12 09.27
HME [6] 32.35 14.02 29.53 12.36 25.13 10.22
HCRN [16] 32.48 16.98 32.01 14.97 25.67 12.23
HGA [15] 31.40 15.36 29.40 13.19 28.23 13.84
MIST [7] 60.51 43.67 59.89 38.05 53.36 36.29
Zero-shot
BLIP-2 [18] 43.67 23.32 45.88 24.18 46.64 26.48
Video-LLaMA [57] 35.00 11.73 35.16 9.34 29.62 9.68
VideoChat2 [21] 38.14 15.36 38.91 15.93 40.83 15.99
Table 1: MCQA Results on our dataset. A: answer; E: explanation.
4.1 Benchmarking on MCQA
Performance metric.
Accuracy is used as the performance metric.
Baselines.
Following prior work [47], we benchmark the performance of the following models on our causal action QA dataset: BlindQA [2]; EVQA [2]; CoMem [8]; HME [6]; HCRN [16]; HGA [15]. Additionally, we also evaluate recent vision-language models like MIST [7], BLIP-2 [18] and multimodal instruction tuning models Video-LLaMA [57], VideoChat2 [21], which obtain solid results on vision-language tasks including VideoQA. We have provided further details on each of these models in Supplementary material.
Evaluation Protocols.
Models are evaluated under the following protocols.
  • Protocol 1. Models’ ability to select the correct Answer is only evaluated. Their ability to select the correct explanation is not taken into consideration. The chance accuracy in this protocol is $1/5$. Results from this protocol are presented under columns marked as A in Table 1.
  • Protocol 2. Models’ ability to select the correct Answer as well as the correct Explanation is evaluated. If the model can select the correct answer, but not the correct explanation, then the model is considered to have failed. In this protocol, the chance accuracy becomes $1/5\times 1/5=1/25$. Results from this protocol are presented under columns marked as A+E in Table 1.
The performances of the baseline models for both protocols are presented in Table 1. In general, we observe that the performances of most of the baseline approaches are low. Most of the models gain only around 11-12% improvement over the chance accuracy for both protocols. MIST seems to be doing exceptionally well compared to other models. We hypothesize that this could be because MIST being geared for long-form video understanding, is better able to handle long video contexts, which are frequent in our dataset. Nonetheless, overall, we observed the following failure modes: 1) Limited Evidence Consideration: They often focus on a small subset of evidence, neglecting other relevant cues distributed throughout the sequence, which is problematic for datasets with long causal chains like ours. 2) Models failed because they tried to exploit shortcuts like object-noun or action-verb matching in video and language spaces instead of focusing on causal relationship modeling—such shortcuts can lead to wrong predictions on our dataset due to our hard negative mining, correct and incorrect answers would likely contain these nouns/verbs. The most effective way to discriminate on our dataset is by inferring the causal relationship. We have discussed further shortcomings and provided qualitative results in Supplementary material.
Comparing the two protocols, we find that Protocol 2, where the model has to select both the correct answer and the correct explanation is significantly more difficult than selecting just the correct answer for all the models. Comparing splits, we observe that, in line with our intuition, UD seems to be the easiest split across all the models. UD is followed by PS; with UN being the most difficult split. This order of difficulty is intuitive. It can also indicate that having at least some understanding of past events or storylines may aid in reasoning about current events or actions.
Negative type Accuracy (%)
Vanilla Hard 60.51
Causal-Confusion 40.57 ($\downarrow 20\%$)
Table 2: Impact of causally confusing options.
We further conducted an experiment where we tested the best performing model, MIST, on our Causal-Confusion set, where incorrect answer choices have the same objects and actors as the correct answer option, but cause and effect relationships are reversed or altered. We observed a significant drop in performance as shown in Table 2. This could potentially be due to causal relationships being not modeled adequately even by such advanced VideoQA models. Interestingly, we also found the S-BERT similarity score to be above 90% for an action (e.g., Tom is running after Jerry) and its Causal-Confusion version (e.g., Tom is not running after Jerry); while these sentences would be opposite/different in terms of human perception. We believe that Causality might be more overlooked than we think in various fields, not just in computer vision.
Model BLEU-1 BLEU-2 BLEU-3 METEOR ROUGE SPICE CIDEr S-BERT CapsMIX
A E A E A E A E A E A E A E A E
BlindQA [2] 0.2412 0.1912 0.1039 0.0618 0.0415 0.0221 0.1243 0.0891 0.2732 0.1885 0.2956 0.0703 0.1034 0.0403 0.4987 0.4195 2.7646
UATT [51] 0.2661 0.2203 0.1257 0.0805 0.0575 0.0330 0.1409 0.1046 0.3171 0.2293 0.3547 0.2775 0.1197 0.0397 0.5495 0.4767 3.3928
HME [6] 0.2650 0.2072 0.1301 0.0713 0.0554 0.0219 0.1504 0.0914 0.3125 0.2273 0.3184 0.0826 0.0936 0.0639 0.5328 0.4737 3.0975
HGA [15] 0.2891 0.2263 0.1588 0.0897 0.0789 0.0323 0.1856 0.1130 0.3388 0.2399 0.3980 0.2426 0.2782 0.0573 0.6026 0.4561 3.7872
BlindGPT-2 0.3770 0.3165 0.2530 0.1749 0.1632 0.1091 0.2452 0.1761 0.3858 0.2966 0.4305 0.3529 1.2482 0.7690 0.6755 0.6271 6.6006
VisionGPT-2 0.3878 0.3095 0.2605 0.1727 0.1738 0.1091 0.2560 0.1756 0.3934 0.2941 0.4498 0.3385 1.3725 0.7539 0.6760 0.6350 6.7582
BLIP-2 [18] 0.1381 0.0815 0.0451 0.0256 0.0167 0.0059 0.0664 0.0480 0.1618 0.1312 0.0530 0.0422 0.2279 0.1046 0.3837 0.3614 1.8931
Video-LLaMA [57] 0.1241 0.1181 0.0419 0.0344 0.0115 0.0098 0.1477 0.1163 0.1719 0.1435 0.2055 0.1383 0.0836 0.0430 0.5734 0.4834 2.4464
VideoChat2 [21] 0.2353 0.2116 0.0823 0.0776 0.0250 0.0253 0.1769 0.1295 0.2667 0.2168 0.3264 0.2547 0.3980 0.2910 0.6445 0.5908 3.9524
(a) (a)
Model BLEU-1 BLEU-2 BLEU-3 METEOR ROUGE SPICE CIDEr S-BERT CapsMIX
A E A E A E A E A E A E A E A E
BlindQA [2] 0.2193 0.1795 0.0772 0.0468 0.0326 0.0149 0.1122 0.0813 0.2501 0.1767 0.2960 0.1147 0.1099 0.0434 0.5129 0.4260 2.6935
UATT [51] 0.2693 0.1947 0.1257 0.0581 0.0440 0.0146 0.1466 0.0847 0.3226 0.2107 0.3255 0.2012 0.1444 0.0646 0.5364 0.4829 3.2260
HME [6] 0.2475 0.1830 0.1031 0.0549 0.0363 0.0154 0.1417 0.0845 0.2820 0.2173 0.2933 0.2549 0.0831 0.0655 0.5165 0.4929 3.0719
HGA [15] 0.2586 0.1842 0.1085 0.0517 0.0365 0.0148 0.1433 0.0932 0.2909 0.1806 0.2908 0.2128 0.0877 0.0266 0.5231 0.4078 2.9111
(b) (b)
Model BLEU-1 BLEU-2 BLEU-3 METEOR ROUGE SPICE CIDEr S-BERT CapsMIX
A E A E A E A E A E A E A E A E
BlindQA [2] 0.2312 0.1847 0.0869 0.0550 0.0319 0.0198 0.1263 0.0930 0.2519 0.1806 0.2758 0.2447 0.0667 0.0265 0.4844 0.4153 2.7747
UATT [51] 0.2659 0.1936 0.1180 0.0645 0.0548 0.0223 0.1423 0.0868 0.3025 0.2200 0.3449 0.2595 0.0940 0.0713 0.5447 0.4750 3.2601
HME [6] 0.2640 0.1930 0.1238 0.0644 0.0556 0.0239 0.1457 0.0868 0.2998 0.2155 0.3115 0.1623 0.0944 0.0698 0.5341 0.4761 3.1207
HGA [15] 0.2674 0.2293 0.1337 0.0823 0.0620 0.0308 0.1703 0.1123 0.3186 0.2363 0.3734 0.2558 0.1972 0.0477 0.5765 0.4692 3.5628
(c) (c)
Table 3: OEAG Results on our dataset. (a) UD split; (b) PS split; (c) UN split.
4.2 Benchmarking on OEAG
CapsMIX Performance metric.
We measure the performance of open-ended generation in terms of BLEU-1,2,3 [33], METEOR [3], ROUGE [23], SPICE [1], CIDEr [44] and Sentence-BERT [36] scores by comparing with the ground truth answers as done by captioning and QA literature. However, we note that with such a wide range of metrics, it is difficult to get a comprehensive insight into models’ performances and compare them. To address that, we introduce a comprehensive metric, termed Caps-MIX (Captioning Metrics Integration eXpert), which integrates all the previously mentioned scores after normalizing them to their theoretical best values. This not only makes it more accessible to compare models based on just a single number but also combines the characteristics of individual metrics, each of which measures performance from a unique perspective. We note that the WUPS [27] score was originally developed for measuring the performance of single-word answers, & may not be an ideal metric for our dataset, which contains longer, detailed answers. Thus, we refrain from using it.
Baselines.
Following prior work [47], we benchmarked the performance of EVQA [2]; UATT [51]; HME [6]; HGA [15] on open-ended answer generation task. Additionally, we also report the zero-shot performance of recent multimodal video understanding models BLIP-2 [18], Video-LLaMA [57], VideoChat2 [21].
We evaluated the baselines for generating answers and explanations. The performances of various models are presented in Table 3. Comparing the BLEU scores, we can see that the performances degrade as we move from BLEU-1 to BLEU-3, indicating that the models are struggling with generating longer answers. We confirmed this qualitatively. Comparing split-type wise, we found that as in the case of MCQA, UD was the easiest split. Overall, we conclude that models struggle quite a bit with open-ended generation.
Inspired by the success of large language models (LLMs), we experimented with leveraging GPT-2 [35], a publicly available LLM, as our natural language answer generator. We found that a pre-trained version out of the box did not work well—it generated mostly random, unrelated words. However, upon simple training on the train set, it performed significantly better than the baselines we considered as shown in Table 3(a) (BlindGPT-2). However, it is unclear if these large language models have “seen" Tom and Jerry scripts during their pretraining stage. If so, then pretraining on relevant scripts followed by finetuning on our dataset could potentially be a reason for GPT-2’s good performance. Although this is less likely. What is more likely to be the reason behind this good performance is the language modeling/generating capability of GPT-2.
In the next step, we integrated visual information into LLM. Taking inspiration from [31], we learn a projector network to align the visual features with GPT-2 representations (VisualGPT-2 in Table 3(a)). We observed an improvement in the performance. However, we believe that this might not be a very efficient way to integrate visual information with LLMs. We expect the performance to boost considerably through better and more sophisticated joint modeling vision and language. Overall, we believe shallower networks might not have the capacity to do inference over longer, complex causal chains, and as such it might not be the best option to invest future efforts into; LLMs have the potential to excel at causal reasoning; this view is also supported in a concurrent survey [25].
Train Data MCQA OEAG
BlindQA HGA MIST HGA
NextQA 32.73 41.23 55.44 1.44
+ Ours 32.91 41.44 55.96 1.48
Table 4: Our dataset improves performance on existing real-world dataset. For OEAG, CapsMIX score is reported.
4.3 Does our dataset help with real-world cases?
To observe our dataset’s direct and immediate impact, we combine our training set with NextQA’s training set in this experiment. Then, we note the performance of models on NextQA test set. For comparison, we also note the performance of models when our dataset is not leveraged/incorporated. Despite being a synthetic/cartoon dataset, our dataset boosted the performance on a real-world dataset on both MCQA and OEAG (Table 4). We observed improvement in identifying the correct causes by breaking the reliance on shortcuts and focus on causal effects; a more comprehensive analysis of the situation/interaction, rather than jumping to a conclusion. Qualitative results provided in Supplementary material. What is more, improvements were not limited to ‘causal-why’ questions but also extended across other question types. We hypothesize that this can attributed to the wider range of reasoning involved in our dataset. To the best of our knowledge, this is the first time a synthetic VideoQA dataset has shown immediate improvement on a real-world dataset. However, we noticed a slight drop in the performance on location-type questions. Thus, we believe that it might be better to: 1) transfer reasoning skills acquired from challenging synthetic datasets like ours; 2) leverage our synthetic dataset to inform the model designing process, as it better reflects the challenges VideoQA models may face in the real world, such as longer causal chains and frequent scene changes in visual streams. Nonetheless, we do not recommend naively transferring or deploying whole models/weights trained on our dataset to real-world scenarios or applications.
5 Conclusion
We introduced CausalChaos!, a challenging dataset for causal action question-answering tasks. Based on the timeless classic of the Tom and Jerry cartoon, it is richly annotated with critical thinking questions demanding a wide spectrum of reasoning from Video QA models. Questions are supported with multi-level answers and further explanations, all-encompassing video context. We further provide a novel CausalConfusion test set, which can challenge the causal relationship modeling capabilities of Video QA models. Through experimentation, we found that while existing models work well on causal action QA tasks, there is a lot of room to improve, especially, in terms of causal relationship modeling, and long and detailed open-ended answer generation. While LLMs offer a promising start, sophisticated integration of visual information with LLMs or joint modeling vision and language would be crucial. We hope that our challenging dataset will foster such developments. We will release our dataset and codes to help future efforts in this direction. Lastly, we showed that it brings improvement to real-world datasets. We believe our dataset, derived from cartoons, should inform the model designing process, as it better reflects the challenges Video-QA models may face in the real world, such as longer causal chains and frequent scene changes in visual streams. However, we do not suggest naively transferring or deploying whole models trained on our dataset to real-world scenarios.
Acknowledgements.
This research/project is supported by the National Research Foundation, Singapore, under its NRF Fellowship (Award# NRF-NRFF14-2022-0001).

### v2

@@
-Ting En Lam1,2, Yuhan Chen1,3, Elston Tan1,4, Eric Peh1, Ruirui Chen1, Paritosh Parmar1, Basura Fernando1
+Yuhan Chen1,3, Elston Tan1,4, Basura Fernando1 1
 Abstract
-Causal video question answering (QA) has garnered increasing interest, yet existing datasets often lack depth in causal reasoning analysis.
+Causal video question answering (QA) has garnered increasing interest, yet existing datasets often lack depth in causal reasoning.
 To address this gap, we capitalize on the unique properties of cartoons and construct CausalChaos!, a novel, challenging causal Why-QA dataset built upon the iconic “Tom and Jerry" cartoon series.
-With thoughtful questions and multi-level answers, our dataset contains much longer causal chains embedded in dynamic interactions and visuals, at the same time principles of animation allows animators to create well-defined, unambiguous causal relationships.
-These factors allow models to solve more challenging, yet well-defined causal relationships.
-We also introduce hard negative mining, including CausalConfusion version.
+Cartoons use the principles of animation that allow animators to create expressive, unambiguous causal relationships between events to form a coherent storyline.
+Utilizing these properties, along with thought-provoking questions and multi-level answers (answer and detailed causal explanation), our questions involve causal chains that interconnect multiple dynamic interactions between characters and visual scenes.
+These factors demand models to solve more challenging, yet well-defined causal relationships.
+We also introduce hard incorrect answer mining, including a causally confusing version that is even more challenging.
 While models perform well, there is much room for improvement, especially, on open-ended answers.
@@
 Along with the other complementary datasets, our new challenging dataset will pave the way for these developments in the field.
-We will release our dataset, codes, and models to help future efforts in this domain.
+1 Introduction
+Understanding the intricate motivations behind human actions is paramount in developing sophisticated systems capable of nuanced behavior analysis.
+In real-world scenarios, actions are shaped by a multitude of factors, including personal experiences, emotions, social contexts, and cultural backgrounds.
+This complexity necessitates a comprehensive approach to unraveling the "why" behind actions, fostering empathy, effective communication, and robust decision-making.
+Causal video question answering aims to decipher the answers behind characters’ actions.
+Despite the growing interest in causal video QA, existing datasets often fall short, requiring only: 1) surface-level understanding; or 2) involve more of simple word substitution in the QA pairs, rather than causal reasoning (e.g., “Q.
+Why are the cars on the street not moving?
+A.
+Parked.").
+Recognizing this gap, we embark on the development of a rigorous and challenging causal Video-QA dataset.
+Our goal is to provide a high-quality resource that rigorously evaluates and advances causal video QA models.
 Figure 1: (Left) Examples of causal questions about characters’ actions from our CausalChaos!
@@
 Please view in Adobe Reader to play the embedded videos for better explanation.
-(Right, top) Illustration of causal chain, scene changes.
+(Middle) Illustration of causal chain, scene changes.
 Linking multiple clues/cues embedded in different scenes to solve causal relationships pose a challenge for VideoQA models.
-(Right, bottom) Being a synthetic medium allows animators to leverage Principles of Animation to stylize the visuals & motions to disentangle/highlight key content of the scene to create well-defined/unambiguous and effectively communicated cause-and-effect relationships.
+(Right) Animators leverage Principles of Animation to stylize the visuals & motions to disentangle/highlight key content of the scene to create well-defined/unambiguous and effectively communicated cause-and-effect relationships.
 The interplay of these factors allow models to focus on solving complex, yet, well-defined, unambiguous causal relationships.
-1 Introduction
-Causal video question answering has been gaining increasing interest [52, 28, 47, 17].
-However, our analysis, as presented in Figs.1, 2, and in § 2, indicates that existing causal QA datasets may often require surface-level understanding or involve more of a simple word substitution in the QA pairs, rather than causal reasoning (e.g., “Q.
-Why are the cars on the street not moving?
-A.
-Parked.").
-Consequently, these datasets do not demand true causal reasoning analysis.
-This gap leaves a dearth of reliable benchmarks for evaluating the causal reasoning capabilities of video understanding models.
-Hence, in this study, we concentrate on developing a rigorous and challenging causal Video-QA dataset.
-We aim to create a high-quality resource that can effectively test and facilitate the advancement of sophisticated causal models.
 Drawing on the established benefits of cartoons with cognitive processes among children [38, 39, 29, 53, 40], we leverage the renowned series [Uncaptioned image] to create a novel and demanding causal Why-QA dataset called CausalChaos!.
@@
 are as follows:
-  • As illustrated in Figs.1, 2, our emphasis lies in crafting thoughtful questions paired with multi-layered answers, aiming to cultivate a deeper and more comprehensive grasp of causal relationships.
+  • As demonstrated in Figure 1, 3, we formulate thought-provoking questions, where the answer (A) and its detailed explanation (E) aim to enhance the model’s understanding of the causal chains of visual events in a given video clip.
+A causal chain is a sequence of events/actions in which each step influences or leads to the next, creating a cause-and-effect relationship.
+It illustrates how one event/action produces a subsequent event/outcome.
 Compared to existing datasets, CausalChaos!
-presents extended chains of causality [^1] A causal chain is a sequence of events/actions in which each step influences or leads to the next, creating a cause-and-effect relationship.
-It illustrates how one event/action produces a subsequent event/outcome., as highlighted in Fig.2 and discussed in § 2.
-  • Video clips in our dataset feature frequent scene and shot changes [^2] A scene change or shot change refers to the transition from one visual setting or perspective to another.
-It typically involves shifting focus to a new location, action, or character within the storyline, often signaled by techniques such as cuts & fades..
-The links or intermediary causes within the causal chains are often dispersed across various scenes.
+presents longer causal chains, as highlighted in Figure 3 and discussed in section 2.
+  • Video clips in our dataset feature frequent scene and shot changes.
+Here, a scene change or shot change refers to the transition from one visual setting or perspective to another.
+It typically involves shifting focus to a new location, action, or character within the storyline.
+The links and causes of the causal chains are often dispersed across various scenes.
 Consequently, models are challenged to exert greater cognitive effort in connecting multiple events (scenes) and identifying intermediate causes to comprehend the “why" questions related to the clips.
-  • Despite the complexity and length of the causal chains, they are distinctly delineated, devoid of ambiguity, and communicated effectively using principles of animation [42] such as staging and exaggeration.
-This deliberate design allows models to concentrate on deciphering causal relationships.
+  • Despite the complexity and length of the causal chains, they are distinctly delineated, unambiguous, and effectively communicated using principles of animation [42] like staging and exaggeration.
+This deliberate design allows models to focus on deciphering causal relationships.
   • CausalChaos!
 introduces an added layer of complexity by necessitating the modeling of actions at various levels of granularity—ranging from sweeping, large-scale movements to nuanced, subtle actions, such as interpreting emotional cues through facial expressions.
-  • Our dataset demands a diverse range of reasoning skills, encompassing deductive, spatial, emotional reasoning, and more, as outlined in Fig.2 and discussed in §3.
-We evaluated various state-of-the-art Video-QA models including recent multimodal instruction tuning models and found that our dataset remains the most challenging causal QA dataset.
-Evaluating on our dataset, helped us reveal that models generally jumped to conclusions based on partial evidence, instead of taking into consideration full set of evidence; models failed to engage in true causal reasoning, trying to select answers based on shortcuts like object/action-noun/verb matching.
-Based on this, we identified more advanced/explicit causal relationship modeling and joint modeling vision and language as the immediate areas for future efforts to focus upon.
-Further, we show that similar to how cartoons help children better connect cause and effects, they can help Video-QA models as well.
+  • Our dataset demands a diverse range of reasoning skills, encompassing deductive, spatial, emotional reasoning, and more, as outlined in Figure 3 and discussed in section 3.
+Upon evaluating various state-of-the-art VideoQA models including the recent multimodal instruction tuning models, we found that our dataset remains one of the most challenging causal QA dataset.
+Particularly, we observed that models often: 1) jumped to conclusions based on partial evidence, rather than considering the full set of evidence; 2) failed to engage in true causal reasoning, opting instead for shortcuts like object/action-noun/verb matching to select answers.
+Based on this, we identified more advanced/explicit causal relationship modeling and jointly modeling vision and language as the immediate areas for future efforts to focus upon.
+Further, we show that similar to how cartoons help children better connect cause and effects, they can help VideoQA models as well.
 We incorporated our dataset with the NextQA [47], a real-world dataset and found that it brings some improvements in why questions.
 What is more, incorporating our dataset brings improvement on non-Why questions as well.
-Overall, we believe that our CausalChaos!
-dataset contributes challenges that span the entire Video-QA pipeline, from deciphering intricate videos to processing complex questions and discerning between nuanced answers, stimulating research in many areas like video processing, causal reasoning, language modeling and joint modeling.
+Overall, we believe our CausalChaos!
+dataset presents challenges spanning the entire VideoQA pipeline, from deciphering intricate videos to processing complex questions and discerning nuanced answers, stimulating research in video processing, causal reasoning, language modeling, and joint modeling.
 Along with the other complementary datasets, such as [47, 17, 19, 21, 46], our new challenging dataset will pave the way for these developments in the field.
-We advocate for leveraging our dataset sourced from cartoons to inform the model design process.
-Our dataset better reflects the challenges Video-QA models may face in the real world, such as longer causal chains and frequent scene transitions in visual streams.
-Nevertheless, we caution against blindly transferring or deploying entire models/weights trained on our dataset to real-world applications.
 Dataset
-Multilevel
-Answers
-Causal
-Chain Len
-Scene
-Reasoning
-Spectrum
-Multigranular
-Actions
-NextQA [47] 1 1 narrow limited
-CausalVidQA [17] 1 1 narrow limited
-IntentQA [19] 1 1 narrow limited
-Ours 2.7 4 wide ubiquitous
+MA
+CCL
+NOS
+RS
+MGA
+NextQA 1.1 1 narrow limited
+CausalVid 0.9 1 narrow limited
+IntentQA 1.3 1 narrow limited
+CausalChaos! 1.9 4 wide abundant
 […]
@@
 and existing causal videoQA datasets.
+MA-multilevel answers; CCL-causal chain length; NOS-no.
+of scenes; RS-reasoning spectrum; MGA-multigranular actions.
 (c) Qualitative comparison between CausalChaos!
-and NextQA datasets.
+and NextQA (Why-QA) datasets.
 CausalChaos!
 Answers and Explanations give detailed information regarding the actual cause-and-effect relationships, motivations, emotions covering wide range of reasoning types.
-Note that, here, with the help of human annotators, we have temporally cropped videos to retain only the relevant parts from NextQA dataset videos; otherwise, raw videos are longer resulting in unintended problem of temporal localization for models.
+Note that, we have temporally cropped videos to retain only the relevant parts from NextQA dataset videos; otherwise, raw videos are longer resulting in unintended problem of temporal localization for models.
 2 Related Work
-VideoQA Datasets and Benchmarks.
-To drive the progress in VideoQA, researchers have developed various datasets with distinct focuses and contributions—see the survey [59].
-In the following, we mainly compare and contrast our work from prior art, with a focus on causal QA literature.
-We provide a detailed discussion on Video QA datasets in the Supplementary material.
-Early VideoQA datasets
-such as MovieFIB [26], MSVD-QA [48], MSRVTT-QA [48], YouTube2Text [10], open-ended QA [58], Zeng et al.
-[56], and Video Context-QA [61] relied on video captions or descriptions to automatically generate questions and corresponding answers, primarily focusing on object and action recognition.
-However, they lacked depth in understanding complex interactions and causal relationships within videos.
-TGIF-QA [13, 14] and ActivityNet-QA [54], while covering short and long web videos respectively, still fell short in capturing comprehensive reasoning, particularly in object interactions and causal relationships.
-Social-IQ [55] addressed human social behaviour in videos through multimodal cues but lacked broader context and visual reasoning capabilities.
-AGQA [9] focused on spatio-temporal scene understanding, but did not cover explanatory "why" questions.
-STAR [46], a situation reasoning dataset, extended to prediction and feasibility questions but also lacked explanatory inquiries.
-Causal QA datasets.
+To drive the progress in VideoQA, researchers have developed various datasets [26, 48, 10, 58, 56, 61, 13, 14, 54, 55, 9, 46] with distinct focuses and contributions—see the survey [59].
+In the following, we mainly compare our work with prior art focusing on causal QA literature.
+We provide a detailed discussion on various VideoQA datasets and models in the Appendix.
 Our work is inspired from a few datasets: CLEVRER [52, 27], NextQA [47], CausalVidQA [17], IntentQA [19] that cover causal reasoning.
-However, they have several limitations that our dataset aims to fill:
-  • Some are limited by their scope, e.g., [52, 28] only consider collision events of simple inanimate objects such as cuboids; as a result, lack: actors, their characteristics such as emotions, intentions, lack diverse language descriptions.
-  • Some datasets (e.g., [47, 17]) do not contain any precise temporal annotations, which inadvertently involves temporal localization & may create a false sense of difficulty, while the questions may not be challenging.
-We alleviate this problem by providing precise temporal annotations and focusing on designing truely complex and challenging QA.
-  • Some datasets lack complexity in questions [47, 17, 19].
-Based on human studies, we found that a majority of causal QA in these datasets were flagged as not requiring any notable reasoning (Figure 2).
+However, they have limitations that our dataset aims to fill:
+  • Limitation in scope: e.g., [52, 27] only consider collision events of simple inanimate objects such as cuboids; as a result, lack: actors, their characteristics such as emotions, intentions.
+Our dataset has actors whose actions are shaped by emotions, intentions, context, etc.
+  • Lack of precise temporal annotations (e.g., [47, 17]) inadvertently involves temporal localization & may create a false sense of difficulty, while the questions may not be challenging.
+We alleviate this problem by providing precise temporal annotations and focusing on designing truly complex and challenging QA.
+  • Lack of complexity in questions (e.g.[47, 17, 19]).
+Based on human studies, we found that a majority of causal QA in these datasets were flagged as not requiring any notable reasoning (Figure 3(a)).
 We also attempted to quantify the complexity in reasoning by computing the lengths of causal chains involved in their QAs vs.
-ours using GPT-4 [32].
-Causal chains in [47, 17, 19] have an average length of 1.
-In comparison, our dataset has much longer causal chains—average length 2.7—posing a challenge for models in connecting multiple events or cues.
-Our dataset also demands wide spectrum of types of reasoning (2(a)).
-  • Scenes in [52, 28, 47, 17, 19] are less dynamic.
-For example, they mostly involve a single scene.
-In comparison, our dataset, on average, has about 4 scene changes.
-Rapid scene changes & dynamic interactions pose a challenge to VideoQA models when modeling cause-&-effect relationships over longer causal chains by linking context and cues embedded in different scenes (more details in subsection 3.2).
-While related challenges have been acknowledged in other computer vision problems [34], they remain unexplored in the realm of VideoQA.
-  • Current datasets do not emphasize on hard incorrect option selection.
-As such, models may not require to do causal reasoning, but rather they can exploit shortcuts like object/action-noun/verb matching in vision-language spaces to correctly answer.
-Thus, we develop hard negative selection strategies.
+ours using GPT-4o [32].
+Causal chains in [47, 17, 19] are shorter (Figure 3(b).
+In comparison, our dataset has longer causal chains—average length 1.9—posing a challenge for models in connecting multiple events or cues.
+Our dataset also demands wide spectrum of types of reasoning (3(a)).
+  • Scenes in [52, 27, 47, 17, 19] are less dynamic, mostly involving a single scene.
+In comparison, our dataset averages about four scenes.
+Rapid scene changes & dynamic interactions challenge VideoQA models to link context & cues across different scenes, modeling cause-and-effect relationships over longer causal chains (more details in subsection 3.2).
+While similar challenges have been acknowledged in other computer vision problems [34], they remain unexplored in VideoQA.
+  • Lack of hard negatives.
+Current datasets may not emphasize on including hard incorrect options.
+Due to this, models may not be required to do causal reasoning, but rather they can exploit shortcuts like object/action-noun/verb matching in vision-language spaces to select correct answer.
+To address this limitation, we develop hard negative selection strategies.
+  • No hierarchical answers with different level of explanation.
 These datasets contain a single-level answer to ‘Why’-questions.
-On the other hand, we richly annotate our dataset to provide two-level answers to ‘Why’-questions.
+On the other hand, we richly annotate our dataset to provide two-level answers to ‘Why’-questions—1) direct/immediate cause; 2) deeper explanation.
 Quality of answers is further enhanced, and are more informative as a consequence of the more complex questions.
-VideoQA Tasks/Models.
-VideoQA generally uses two protocols: 1) multichoice QA (MCQA) & 2) open-ended answer generation (OEAG).
-MCQA involves choosing the correct option from the multiple provided options.
-MCQA models [2, 8, 6, 16, 15, 7] fuse Video, question, & answer representations & process them using a multi-layer perceptron, which ultimately predicts the correct answer.
-On the other hand, in much more challenging OEAG, VideoQA models generate answers in natural language [2, 51, 6, 15].
-This is typically accomplished by decoding the fused video & question representations into a sequence of words (answer).
-Recently, vision-language and multimodal instruction tuning models [7, 18, 57, 21] have shown strong performance on video understanding tasks.
-We evaluate & report the performance of all these models on our dataset.
-Further discussion on individual models can be found in Supplementary material.
 3 CausalChaos! Dataset
-In this section, we detail the construction of our challenging VideoQA dataset focused on causal reasoning.
-We provide insights into the video source, annotation process, and rigorous quality checks employed to ensure detailed annotations of high quality.
-Further, we discuss the unique attributes of our dataset that make it a valuable resource for the causal action/VideoQA task.
+This section details the construction of our VideoQA dataset designed to challenge causal reasoning, covering its video source, annotation process, & quality checks to ensure high-quality annotations.
+We then discuss the unique attributes that make our dataset valuable for causal VideoQA tasks.
 3.1 Dataset Construction
 Video source.
-To exclusively focus on visual information for reasoning, we selected the timeless silent cartoon series, ‘Tom & Jerry’ (1940), spanning over 6 seasons and encompassing a total of 161 episodes.
-Silent cartoons that do not contain modalities other than vision (such as dialogues) can foster the development of visual reasoning capabilities in children.
-Similarly, we hypothesize that using silent cartoons for VideoQA can allow for a focus on visual reasoning.
-The Tom & Jerry series is also an ideal source for constructing a causal reasoning dataset, as it provides abundant segments showcasing diverse cause-&-effect relationships.
-A concurrent work [22] on causal image generation has also leveraged Tom & Jerry cartoon.
+To focus on visual reasoning, we selected the timeless silent cartoon series, “Tom & Jerry" (1940) [^1] We only provide the annotations; Videos can be obtained from: https://www.warnerbros.com/tv/tom-and-jerry-1965-volume-1., spanning over six seasons and 161 episodes.
+Silent cartoons, which lack other modalities like dialogue, foster visual reasoning in children.
+Similarly, we hypothesize that using silent cartoons for VideoQA will enhance visual reasoning.
+The Tom & Jerry series is ideal for a causal reasoning dataset, offering abundant segments with diverse cause-and-effect relationships.
+Concurrent work [22] on causal image generation also leverages Tom & Jerry.
 Annotations.
-We collect the following annotations.
-$\bullet$ Questions.
-The questions are crafted to capture the why or reasoning behind the actions of characters in Tom & Jerry cartoons.
-To cultivate comprehensive and deeper video understanding capabilities in models, we formulated thoughtful causal questions where cause & effect are connected by longer causal chains.
-Questions also extend beyond explicit visual cues, encompassing gestures & expressions, to delve into the characters’ underlying intentions and goals.
-The annotators would generate questions while watching the video for the first time to ensure that they did not have prior knowledge of the scene, which mimics how models would assess unseen clips.
-Subsequently, they would rewatch the videos multiple times to generate more critical thinking questions.
-While generating questions, annotators watch the videos without audio to ensure that the dataset does not contain audio cues to focus on visual reasoning.
+Each dataset sample includes the following annotations: {start frame, end frame, Question, Answer, Explanation}. Details in the following.
+$\bullet$ Questions are crafted to capture the why or reasoning behind the actions of characters in Tom & Jerry cartoons.
+To cultivate comprehensive & deeper video understanding capabilities in models, we formulated thought-provoking causal questions where cause & effect are connected by longer causal chains.
+Questions also extend beyond explicit visual cues, encompassing gestures & expressions, to delve into the characters’ underlying intentions & goals.
+Annotators generated questions while watching the video for the first time to mimic how models assess unseen clips.
+They then rewatched the videos multiple times to create more critical thinking questions.
+To focus on visual reasoning, annotators watched the videos without audio, ensuring no audio cues were included in the dataset.
 $\bullet$ Multi-level Answers.
-The first-level or the Primary answer, represents a shallower, literal or direct cause or form of response.
+The first-level or the Primary answer, represents a literal or direct cause or form of response.
 It is accompanied by a deeper form of Explanation, which considers the broader context of the scene, the thoughts, feelings, intentions of characters, & their actions.
@@
 It includes reasons & additional details to comprehensively address the question.
-Examples shown in Figure 1 & Supplementary material.
-$\bullet$ Temporal annotations.
-For each QAE set the Start &End frame numbers of the video are recorded such that the entire scene including frames that offer context & details supporting the reasons behind the actions is covered.
-In summary, each dataset sample is annotated with: {start frame, end frame, Question, Answer, Explanation}.
+Examples shown in Figure 1 & Appendix.
+$\bullet$ Temporal annotations. For each QAE set, the Start and End frame numbers are recorded to cover the entire scene, including contextual frames that support the reasons behind the actions.
 Quality Check.
-To ensure the quality and reliability of the dataset, we implemented a two-stage quality check process.
-In the first stage, quality checkers conducted thorough assessments of episodes they did not annotate.
-Subsequently, in the second stage, quality checkers reviewed episodes for which they did not perform the first-stage quality check.
-Multiple quality checkers were involved in the process, carefully reviewing and verifying curated question and answer sets for logical fallacies, timestamp inconsistencies, grammatical errors, and spelling mistakes.
-The process also involved checking and excluding annotations containing audio or text aspects that the model cannot comprehend.
-To prevent direct overwriting, any edits were flagged in a different font color for further discussion.
-The inconsistencies or errors that were flagged out were resolved through discussions and consensus among the checkers and respective annotators.
-In these discussions, the team referred to the episode synopses available on Tom and Jerry Wiki [43], an online Tom and Jerry encyclopedia.
-This external source served as a third-party opinion and facilitated fact-checking.
-By incorporating multiple perspectives, the discussions helped to provide a deeper level of analysis to generate a more challenging and comprehensive dataset.
+To ensure dataset quality and reliability, we implemented a two-stage quality check process.
+First, quality checkers assessed episodes they did not annotate.
+Then, they reviewed episodes they did not check in the first stage.
+Multiple quality checkers carefully reviewed and verified curated question and answer sets for logical fallacies, timestamp inconsistencies, grammatical errors, and spelling mistakes.
+They also excluded annotations containing audio or text aspects that the model cannot comprehend.
+To prevent direct overwriting, edits were flagged in a different font color for further discussion.
+Flagged inconsistencies or errors were resolved through discussions and consensus among checkers and annotators.
+The team referred to episode synopses from Tom and Jerry Wiki [43], an online encyclopedia, as a third-party opinion and for fact-checking.
 Dataset Statistics.
-Annotators meticulously watched all 161 cartoon episodes spanning 6 seasons, generating a total of 4945 richly annotated Question-Answer-Explanation sets for the dataset.
-The inclusion of a substantial number of Question-Answer-Explanation sets across all seasons, covering a wide range of scenes, characters, and events, underscores the comprehensive nature of our dataset.
-This diversity establishes a robust foundation for understanding and answering questions related to the cartoon series.
-Word clouds corresponding to answers and explanations from our dataset are presented in Supplementary material.
-3.2 Unique Characteristics of our dataset
+Annotators reviewed all 161 episodes across 6 seasons, generating 4,945 detailed Question-Answer-Explanation sets.
+This comprehensive dataset covers various scenes, characters, and events, providing a solid foundation for understanding and answering questions about the series.
+Additional statistics and word clouds for answers and explanations are available in the Appendix.
+Figure 4: Grounded in diverse visual information.
+Figure 5: Traditional temporal modeling vs Dynamic scene linking. Notice the abrupt scene change, which causes disruption in visual flow, resulting in large amplitude and widespread optical flow.
+3.2 Unique Characteristics of our CausalChaos! Dataset
 In the following, we delve into the untapped unique properties of our dataset, which both complement existing video understanding datasets and present novel challenges to VideoQA, particularly from a causal reasoning perspective.
   1. 1.
-    Multi-level answers.
-VideoQA datasets generally provide single-level answers.
-Single-level answers might contain less information.
-Multi-level answers are valuable because they offer a richer and more nuanced perspective, which leads to deeper insights—considering not only the immediate circumstances but also underlying psychological, social, and personal factors that contributed to characters’ decisions.
-By employing multi-level answers alongside thoughtful questions, we can cultivate deeper reasoning and analysis of characters’ actions, enabling the training and evaluation of models on complex question-answer pairs that demand a comprehensive understanding of longer causal chains.
-Further elaborated discussion on the importance of multi-level answers to ‘Why’-questions included in Supplementary material.
+    Multi-level answers offer a richer and more nuanced perspective, which leads to deeper insights—considering not only the immediate circumstances but also underlying psychological, social, and personal factors that contributed to characters’ decisions.
+By employing multi-level answers alongside thought-provoking questions, we can cultivate deeper reasoning and analysis of characters’ actions, enabling the training and evaluation of models on complex question-answer pairs that demand a comprehensive understanding of longer causal chains.
+Further elaborated discussion on the importance of multi-level answers to ‘Why’-questions included in Appendix.
   2. 2.
     Grounded in diverse Visual & Motion information.
-The Tom and Jerry cartoon series is visually rich and dynamic with intricate scenes, actions, and character interactions varying across all episodes.
-We asked annotators to craft questions where the answer is grounded in the video.
-By grounding questions(-answers) in the visual input, the model is required to analyze the video to gain access to a broader range of details and cues/clues from the videos to provide meaningful answers.
-For example, the same question—Why was Jerry laughing?—appears in two different seasons (episodes S03E17, S05E13).
-However, the Explanations are different: 1) for S03E17, the Explanation is “Jerry found it funny that the seal stole the fish from Tom."; 2) while for S05E13, the Explanation is “Jerry was laughing at Tom who fell down the stairs and into the fountain." This concretely shows that VideoQA models need to understand visual information to correctly answer the questions.
-Grounding QAE set in vision is underscored by the presence of unusual situations in cartoons—typically not found in real-world cases.
-For example, Jerry removing Tom’s nose and fitting strawberry in place of Tom’s nose or Lion hushing Jerry.
-Due to this, models cannot answer just based on the correlations/biases in its training data, rather models ought to process and understand the contents thoroughly to answer correctly.
-This is crucial even in with large language models (LLMs)/embeddings, which are also susceptible to such frequency-based biases/hallucinations [20, 24, 60].
-For example, Lion is closely related to killing in embeddings from LLMs; in a qualitative analysis, we found that models may tend to choose an answer containing ‘killing’ Jerry as the motive for Lion-character’s action, while in actuality the Lion-character was trying to quiet Jerry.
+The Tom & Jerry cartoon series is visually rich and dynamic, with intricate scenes, actions, and character interactions varying across episodes.
+We had annotators craft questions where the answer is grounded in the video.
+This grounding compels the model to analyze the video for a broader range of details and clues to provide meaningful answers.
+For example, consider Figure 4(a), same question appears in different episodes, but their answers/context are completely different.
+This demonstrates that it is crucial for VideoQA models to understand visual information to answer correctly on our dataset.
+Multimodal nature of VideoQA task is enhanced by the presence of unusual situations in cartoons—typically not found in real-world datasets as shown in Figure 4(b,c).
+Since the answers are embedded in videos, models cannot rely on correlations/biases in their training data; they must thoroughly process and understand the content to answer correctly.
+This is crucial even with large language models (LLMs) and embeddings, which are also susceptible to frequency-based biases and hallucinations [20, 24, 60].
+For example, Lion is often associated with killing in LLM embeddings.
+In a qualitative analysis, we found that models might choose an answer involving ’killing’ Jerry as the motive for the Lion-character’s action, while the Lion-character was actually trying to quiet Jerry Figure 4(b-left).
 Revealing such biases has been considered in other computer vision problems [4, 12], but is yet to be introduced in VideoQA.
@@
     Focus on causal reasoning in visually dynamic scenes.
-Our dataset focuses on causal reasoning in Tom and Jerry cartoons, which have an abundance of dynamic and rapidly changing scenes, making it challenging for traditional video understanding models to track and comprehend the context effectively.
-To effectively model cause-and-effect relationships and answer causal questions, video understanding models must comprehend the temporal and contextual flow of events, linking them together amid a multitude of rapid scene changes and dynamic interactions.
-Furthermore, characters and objects may appear, disappear, and reappear, potentially limiting the model’s access to only partial observations.
-We term this as Dynamic Scene Linking (DSL).
-DSL is different from temporal modeling in traditional sense, which is typically concerned with modeling temporal evolution of events within a scene (more details in Supplementary material).
-Related challenges have been recognized in other computer vision problems [34, 45], but not yet in VideoQA.
-We hypothesize that humans effectively understand these cartoons by forming a mental ‘world model’ of the scene.
-This mental model aids in bridging gaps between discontinuous scenes and provides a coherent representation of the overall context.
-Incorporating such capabilities into VideoQA has the potential to enhance model performance.
-Notably, current VideoQA datasets often lack coverage of such complex scenarios, unlike our dataset.
+Our dataset focuses on causal reasoning in Tom and Jerry cartoons, characterized by dynamic and rapidly changing scenes.
+Characters and objects may appear, disappear, and reappear, limiting the model’s access to partial observations.
+Video understanding models struggle to track and understand context.
+To answer causal questions, these models must link events to form a causal chain despite rapid scene changes and dynamic interactions.
+We term this Dynamic Scene Linking (DSL), distinct from traditional temporal modeling, which typically focuses on the gradual transitions within a scene as shown in Figure 5 (more details in Appendix).
+Related challenges are noted in other computer vision problems [34, 45] but not yet in VideoQA.
+We hypothesize that humans understand cartoons by forming a mental “world model" of the scene, which helps bridge gaps between discontinuous scenes and provides a coherent overall context.
+Current VideoQA datasets often lack coverage of such complex scenarios, unlike our dataset.
   4. 4.
     More challenging and requiring cognitive effort.
-Understanding causal chains can help explain why things happen and how different factors influence each other.
-Identifying and solving longer causal chains in the process of answering causal (why) questions involve linking together multiple cues/clues, which requires significant cognitive efforts on the part of VideoQA models, especially, when the clues/cues are embedded in different scenes (described above) like in our dataset.
-Annotators were explicitly instructed to craft QAE pairs that require linking together multiple cues under multiple scene changes.
-It is difficult to quantify the complexity of causal questions.
-To the best of our knowledge, there are no existing approaches to accomplish this.
-To this end, we propose using Causal Chain Length as a way to measure the complexity of causal questions.
-Computing lengths of Causal Chains.
-We leverage GPT-4 [32] to compute the lengths of causal chains of 100 randomly chosen QAE pairs from our dataset and existing causal(-Why) VideoQA datasets [47, 17, 19].
-Results are presented & compared in Figure 2.
-We observed that the QAE pairs in our dataset have significantly longer causal chains than existing QA datasets.
-This finding sheds light on the unique challenges posed by our dataset and underscores the importance of developing robust VideoQA models capable of handling complex causal reasoning tasks.
-Examples of causal chains and additional details on computing causal chain lengths provided in Supplementary material.
+We used GPT-4o [31] to compute the lengths of causal chains of 100 randomly chosen QAE pairs from our dataset and existing causal(-Why) VideoQA datasets [47, 17, 19].
+Results are presented in Figure 3.
+We observed that the QAE pairs in our dataset have longer causal chains than existing QA datasets, potentially suggesting our QAs are more complex.
+Identifying and solving longer causal chains in the process of answering causal (why) questions involve linking together multiple cues/clues, which requires significant cognitive efforts on the part of VideoQA models, especially, when the clues/cues are embedded in different scenes (described above), as in our dataset.
   5. 5.
@@
 Overall, causal relationships in our dataset are complex, longer, but at the same time they are unambiguous/well-defined using principles of animation, allowing models to focus on deciphering causal relationships.
-[…]
-    [loop,autoplay,poster=1,width=0.95]3Figs/tor_frs/fr_0015
-    Figure 3: Examples of various types of reasoning required by our dataset. Please zoom in and view in AdobeReader to watch the embedded videos for better explanation.
+    Figure 6: Examples of various types of reasoning required by our dataset. Please zoom in & view in AdobeReader to play the embedded videos. Further types of reasoning visualized in Appendix.
   6. 6.
@@
 It is also possible that a question may involve more than one type of reasoning.
-We conducted human studies to determine the types of reasoning required by our dataset and NextQA and observed that our dataset demands a wider range of reasoning than NextQA (results in 2(a)).
-More information and definitions of types of reasoning and human studies are provided in Supplementary material.
-  7. 7.
-    Character names used instead of general nouns.
-In our dataset, we employ character names rather than generic nouns, enhancing contextual understanding and knowledge representation.
-The utilization of character names necessitates character/actor resolution, encouraging Video QA models to better comprehend question context and character dynamics.
-This contextual understanding results in more precise and contextually fitting answers.
-This approach becomes particularly crucial in scenes with multiple instances of the same type of animals, such as several cats or mice.
-Relying on character names ensures that models are required to distinguish between individual characters and accurately identify who is interacting with whom.
-Failure to discern these interactions may lead to inaccurate responses.
+We conducted human studies to determine the types of reasoning required by our dataset and NextQA and observed that our dataset demands a wider range of reasoning than NextQA (results in 3(a)).
+More information and definitions of types of reasoning and human studies are provided in Appendix.
 3.3 Benchmarks
 Question-types/Tasks.
-Our dataset comprises of following types of questions.
-  1. 1.
-    Multiple Choice Question Answering.
+Our dataset comprises two types of questions.
+The first is Multiple Choice Question Answering (MCQA).
 VideoQA models are provided with the question, the associated video footage, the correct answer, and four incorrect answers; the task for the models is to pick the correct answer.
-  2. 2.
-    Open-ended Answer Generation. The VideoQA models are provided only with the question and the video footage, and the models have to generate the answers in natural language.
+The second is Open-ended Answer Generation (OEAG).
+The models are provided only with the question and the video footage, and the models have to generate the answers in natural language.
 Incorrect answer/option mining.
-In designing MCQA, the answer options should be sufficiently distinct from both the correct answer & each other.
-However, they should also possess enough semantic similarities to challenge reasoning capabilities beyond common sense.
-This way models cannot answer correctly by taking shortcuts like simple object/action matching in vision-language spaces.
-Naive Negative mining or using randomly chosen answers from other questions as incorrect options is unproductive; it renders the correct answer too obvious, so we avoid this strategy.
-Rather, we introduce the following strategies.
+In designing MCQA, answer options should be distinct from the correct answer & each other but semantically similar enough to challenge reasoning beyond commonsense.
+This prevents models from taking shortcuts like simple object/action matching in vision-language spaces.
+Using randomly chosen answers as incorrect options is unproductive and makes the correct answer too obvious, so we avoid this strategy.
+Instead, we introduce these strategies:
   1. 1.
     Vanilla Hard Negative mining.
-A pre-trained Sentence-Bert was adopted in the measure of semantic similarity.
-For each question of interest, we first sample top-10 other questions based on the cosine-similarity measure of the Sentence-Bert embedding.
-From these 10 question candidates, the associated answers are then sampled, and the cosine-similarity measure is once again applied between the correct answer and the answer candidate.
-Using a similarity measure in the equation: $\alpha cossim(Q_{i},Q_{n})\cdot\beta cossim(A_{i},A_{n})$, where $\alpha$ and $\beta$ can be varied to adjust the weights coming from the questions or answers, the top 4 answers are included as the option candidates.
-To prevent candidates from being too similar or rephrasing the form of the answer, the measure heavily penalizes answers with a high similarity $cossim(A_{i},A_{n})$ by setting $\beta$ = 0.001 to reduce the overall score of the candidate.
+A pre-trained Sentence-Bert measures semantic similarity.
+For each question, we sample the top-10 other questions based on cosine similarity of Sentence-Bert embeddings.
+From these, we sample associated answers and again apply cosine similarity between the correct and candidate answers.
+Using the equation $\alpha cossim(Q_{i},Q_{n})\cdot\beta cossim(A_{i},A_{n})$, with adjustable weights $\alpha$ and $\beta$, we select the top 4 answer candidates.
+To avoid overly similar answers, we penalize high similarity scores by setting $\beta$ to 0.001, reducing the candidate’s overall score.
   2. 2.
-    Causal-Confusion Negative mining.
-We leverage the power of large language models to generate confusing incorrect options which have the opposite effect (e.g.
-Tom was hitting Jerry $\rightarrow$ Tom was not hitting Jerry) or the causal agents are swapped (e.g.
-Tom was hitting Jerry $\rightarrow$ Jerry was hitting Tom).
-These causally-opposite incorrect options play an important role in causal action question-answering datasets like ours.
-We hypothesize that the models might not be doing sufficient joint modeling of language and vision.
-As such, Video QA models might be able to answer based on matching actors and objects in questions and answers.
-However, when tested on Causally-Opposite incorrect options, simple question-answer matching might not be enough, and consequently, Video QA models may fail to answer correctly.
-We hypothesize that simple objects and actors might be modeled by VideoQA models, but relationships might not be as they require more dynamic features and as such are harder to model.
-We have provided examples from Causal-Confusion vs.
-Vanilla Hard Negative in the Supplementary material.
+[…]
 Design of Train-Val-Test splits.
-In general, we divide the dataset into a train-val-test split ratio of 70:15:15. We also use the following train/test strategies.
+We typically partition the dataset into 70% for training, and 15% each for validation and testing. Additionally, we employ specific strategies for training and testing.
   1. 1.
-    Seen-episode testing.
-    1. (a)
     Uniformly Distributed Seen-episode testing (UD).
-In this case, dataset samples are randomly divided into train-val-test splits.
-No constraints such as the train samples should be drawn only from certain episodes or only from certain parts of certain episodes are imposed.
-Due to this, it is likely that uniformly distributed chunks/portions of all episodes from all seasons are seen during training.
-This way the VideoQA models can potentially have sparse storylines with them.
-    2. (b)
+Dataset samples are randomly divided into train-val-test splits without constraints on which episodes or parts are included in each split.
+Consequently, uniformly distributed chunks of all episodes from all seasons are likely seen during training, potentially giving VideoQA models sparse storylines.
+  2. 2.
     Consecutive Partially Seen-episode testing (PS).
-In this case, train set is constrained to be drawn from starting 70% video of each episode; val set is constrained to be drawn from the middle 70-85% video of each episode; while the test set is constrained to be drawn from the remaining last 15% of the video of each episode.
-This creates a scenario where the model has partially “seen" episodes during training stage and then does “educated guessing" when answering during testing stage.
-  2. 2.
+The train set is drawn from the first 70% of each episode, the val set from the middle 70-85%, & the test set from the last 15%.
+This setup allows the model to have partially "seen" episodes during training & make "educated guesses" during testing.
+  3. 3.
     Unseen-episode testing (UN).
@@
 4 Experiments
-We benchmark the performance of state-of-the-art VideoQA methods on our newly introduced causal action QA dataset on two tasks: 1) MCQA; 2) OEAG.
-UD PS UN
-Model A A+E A A+E A A+E
-Chance 20.00 04.00 20.00 04.00 20.00 04.00
-BlindQA [2] 29.38 13.07 26.51 11.54 25.13 11.02
-EVQA [2] 29.38 13.48 31.32 13.32 27.82 14.65
-CoMem [8] 32.08 13.88 26.10 09.89 23.12 09.27
-HME [6] 32.35 14.02 29.53 12.36 25.13 10.22
-HCRN [16] 32.48 16.98 32.01 14.97 25.67 12.23
-HGA [15] 31.40 15.36 29.40 13.19 28.23 13.84
-MIST [7] 60.51 43.67 59.89 38.05 53.36 36.29
-Zero-shot
-BLIP-2 [18] 43.67 23.32 45.88 24.18 46.64 26.48
-Video-LLaMA [57] 35.00 11.73 35.16 9.34 29.62 9.68
-VideoChat2 [21] 38.14 15.36 38.91 15.93 40.83 15.99
-Table 1: MCQA Results on our dataset. A: answer; E: explanation.
+We benchmark SOTA VideoQA methods on our new causal action QA dataset for two tasks: 1) MCQA and 2) OEAG. Then, we explore if our dataset aids real-world QA.
 4.1 Benchmarking on MCQA
-Performance metric.
-Accuracy is used as the performance metric.
 Baselines.
-Following prior work [47], we benchmark the performance of the following models on our causal action QA dataset: BlindQA [2]; EVQA [2]; CoMem [8]; HME [6]; HCRN [16]; HGA [15].
-Additionally, we also evaluate recent vision-language models like MIST [7], BLIP-2 [18] and multimodal instruction tuning models Video-LLaMA [57], VideoChat2 [21], which obtain solid results on vision-language tasks including VideoQA.
-We have provided further details on each of these models in Supplementary material.
+Following prior work [47], we benchmark the performance of these models on our causal action QA dataset: BlindQA [2], EVQA [2], CoMem [8], HME [6], HCRN [16], and HGA [15].
+We also evaluate recent vision-language models like MIST [7], BLIP-2 [18], and multimodal instruction tuning models Video-LLaMA [57] and VideoChat2 [21], GPT-4o [31] which excel on vision-language tasks.
+Further details on each model are provided in Appendix.
 Evaluation Protocols.
-Models are evaluated under the following protocols.
+Models are evaluated under these protocols, with Accuracy as the metric:
   • Protocol 1.
-Models’ ability to select the correct Answer is only evaluated.
+Models’ ability to select the correct Answer (A) is only evaluated.
 Their ability to select the correct explanation is not taken into consideration.
@@
   • Protocol 2.
-Models’ ability to select the correct Answer as well as the correct Explanation is evaluated.
+Models’ ability to select the correct Answer (A) as well as the correct Explanation (E) is evaluated.
 If the model can select the correct answer, but not the correct explanation, then the model is considered to have failed.
@@
 Results from this protocol are presented under columns marked as A+E in Table 1.
+UD PS UN
+Model A A+E A A+E A A+E
+Chance 20.00 04.00 20.00 04.00 20.00 04.00
+BlindQA 29.38 13.07 26.51 11.54 25.13 11.02
+EVQA 29.38 13.48 31.32 13.32 27.82 14.65
+CoMem 32.08 13.88 26.10 09.89 23.12 09.27
+HME 32.35 14.02 29.53 12.36 25.13 10.22
+HCRN 32.48 16.98 32.01 14.97 25.67 12.23
+HGA 31.40 15.36 29.40 13.19 28.23 13.84
+MIST 62.22 44.88 62.22 42.86 55.91 37.90
+MIST-CC 63.34 46.80 62.50 43.54 56.18 39.25
+Zero-shot
+BLIP-2 43.67 23.32 45.88 24.18 46.64 26.48
+Video-LLaMA 35.00 11.73 35.16 9.34 29.62 9.68
+VideoChat2 38.14 15.36 38.91 15.93 40.83 15.99
+GPT-4o 63.64 48.17 63.84 49.79 64.82 52.23
+Table 1: MCQA Results on our dataset.We design multitask version of MIST that learns to generate causal chains as an auxiliary task.
+Quantitative results.
 The performances of the baseline models for both protocols are presented in Table 1.
@@
 We hypothesize that this could be because MIST being geared for long-form video understanding, is better able to handle long video contexts, which are frequent in our dataset.
-Nonetheless, overall, we observed the following failure modes: 1) Limited Evidence Consideration: They often focus on a small subset of evidence, neglecting other relevant cues distributed throughout the sequence, which is problematic for datasets with long causal chains like ours.
-2) Models failed because they tried to exploit shortcuts like object-noun or action-verb matching in video and language spaces instead of focusing on causal relationship modeling—such shortcuts can lead to wrong predictions on our dataset due to our hard negative mining, correct and incorrect answers would likely contain these nouns/verbs.
+Overall, we observed the following failure modes for models: 1) Limited Evidence Consideration: Models often focus on a small subset of evidence, neglecting other relevant cues distributed throughout the sequence, which is problematic for datasets with long causal chains like ours.
+2) Models failed because they tried to exploit shortcuts like object-noun or action-verb matching in video-language spaces instead of focusing on causal relationship modeling—such shortcuts can lead to wrong predictions on our dataset due to our hard negative mining, correct & incorrect answers would likely contain these nouns/verbs.
 The most effective way to discriminate on our dataset is by inferring the causal relationship.
-We have discussed further shortcomings and provided qualitative results in Supplementary material.
+We have discussed further shortcomings with qualitative results in Appendix.
+To mitigate these shortcomings, we design a MIST-based multitask model, MIST-CC, which, in addition to doing VideoQA, also learns to generate the causal chain from Video-Question features (groundtruth causal chains are generated by GPT-4 from QAE pairs).
+We found that this simple modification mildly boosts the performance.
+We have provided further details on MIST-CC in Appendix.
+Moving to VLM/MLLMs, we observe that they perform better than traditional models ([2, 15, 16]), except MIST.
+However, GPT-4o released only in late May 2024, outperformed all the models.
+Although, it is closed source model, we hypothesize its vision and language capabilities maybe significantly better than other VLMs.
+It was able to incorporate small details like facial expressions.
+To glean insights into it, we also asked it to give us its reasoning, and found that it analyzes each option individually and then selects the most likely answer.
+This is close to (or at least mimicking) how humans would approach this task.
+We believe this, at least on surface, seems to be going beyond just correlation-based answer picking as in non-VLM/MLLMs.
 Comparing the two protocols, we find that Protocol 2, where the model has to select both the correct answer and the correct explanation is significantly more difficult than selecting just the correct answer for all the models.
-Comparing splits, we observe that, in line with our intuition, UD seems to be the easiest split across all the models.
-UD is followed by PS; with UN being the most difficult split.
-This order of difficulty is intuitive.
-It can also indicate that having at least some understanding of past events or storylines may aid in reasoning about current events or actions.
-Negative type Accuracy (%)
-Vanilla Hard 60.51
-Causal-Confusion 40.57 ($\downarrow 20\%$)
-Table 2: Impact of causally confusing options.
-We further conducted an experiment where we tested the best performing model, MIST, on our Causal-Confusion set, where incorrect answer choices have the same objects and actors as the correct answer option, but cause and effect relationships are reversed or altered.
-We observed a significant drop in performance as shown in Table 2.
+Comparing splits, We observe that UD is the easiest split across all models, followed by PS, with UN being the most difficult.
+This intuitive order suggests that understanding past events or storylines may aid in reasoning about current events or actions.
+Negative type MIST MIST-CC GPT-4o
+Vanilla Hard 62.22 63.34 63.64
+Causal-Confusion 55.80 58.76 54.95
+Table 2: Impact of causal-confusion.
+We further conducted an experiment where we tested the best performing models, MIST, MIST-CC, & GPT-4o, on our Causal-Confusion set, where incorrect answer choices have the same objects and actors as the correct answer option, but cause-and-effect relationships are reversed or altered.
+We observed a significant drop in performance of all models as shown in Table 2.
 This could potentially be due to causal relationships being not modeled adequately even by such advanced VideoQA models.
+Performance of MIST-CC dropped relatively less, potentially because the auxiliary task of generating causal chains may have enhanced the understanding of causal relationships.
 Interestingly, we also found the S-BERT similarity score to be above 90% for an action (e.g., Tom is running after Jerry) and its Causal-Confusion version (e.g., Tom is not running after Jerry); while these sentences would be opposite/different in terms of human perception.
 We believe that Causality might be more overlooked than we think in various fields, not just in computer vision.
-Model BLEU-1 BLEU-2 BLEU-3 METEOR ROUGE SPICE CIDEr S-BERT CapsMIX
-A E A E A E A E A E A E A E A E
-BlindQA [2] 0.2412 0.1912 0.1039 0.0618 0.0415 0.0221 0.1243 0.0891 0.2732 0.1885 0.2956 0.0703 0.1034 0.0403 0.4987 0.4195 2.7646
-UATT [51] 0.2661 0.2203 0.1257 0.0805 0.0575 0.0330 0.1409 0.1046 0.3171 0.2293 0.3547 0.2775 0.1197 0.0397 0.5495 0.4767 3.3928
-HME [6] 0.2650 0.2072 0.1301 0.0713 0.0554 0.0219 0.1504 0.0914 0.3125 0.2273 0.3184 0.0826 0.0936 0.0639 0.5328 0.4737 3.0975
-HGA [15] 0.2891 0.2263 0.1588 0.0897 0.0789 0.0323 0.1856 0.1130 0.3388 0.2399 0.3980 0.2426 0.2782 0.0573 0.6026 0.4561 3.7872
-BlindGPT-2 0.3770 0.3165 0.2530 0.1749 0.1632 0.1091 0.2452 0.1761 0.3858 0.2966 0.4305 0.3529 1.2482 0.7690 0.6755 0.6271 6.6006
-VisionGPT-2 0.3878 0.3095 0.2605 0.1727 0.1738 0.1091 0.2560 0.1756 0.3934 0.2941 0.4498 0.3385 1.3725 0.7539 0.6760 0.6350 6.7582
-BLIP-2 [18] 0.1381 0.0815 0.0451 0.0256 0.0167 0.0059 0.0664 0.0480 0.1618 0.1312 0.0530 0.0422 0.2279 0.1046 0.3837 0.3614 1.8931
-Video-LLaMA [57] 0.1241 0.1181 0.0419 0.0344 0.0115 0.0098 0.1477 0.1163 0.1719 0.1435 0.2055 0.1383 0.0836 0.0430 0.5734 0.4834 2.4464
-VideoChat2 [21] 0.2353 0.2116 0.0823 0.0776 0.0250 0.0253 0.1769 0.1295 0.2667 0.2168 0.3264 0.2547 0.3980 0.2910 0.6445 0.5908 3.9524
-(a) (a)
-Model BLEU-1 BLEU-2 BLEU-3 METEOR ROUGE SPICE CIDEr S-BERT CapsMIX
-A E A E A E A E A E A E A E A E
-BlindQA [2] 0.2193 0.1795 0.0772 0.0468 0.0326 0.0149 0.1122 0.0813 0.2501 0.1767 0.2960 0.1147 0.1099 0.0434 0.5129 0.4260 2.6935
-UATT [51] 0.2693 0.1947 0.1257 0.0581 0.0440 0.0146 0.1466 0.0847 0.3226 0.2107 0.3255 0.2012 0.1444 0.0646 0.5364 0.4829 3.2260
-HME [6] 0.2475 0.1830 0.1031 0.0549 0.0363 0.0154 0.1417 0.0845 0.2820 0.2173 0.2933 0.2549 0.0831 0.0655 0.5165 0.4929 3.0719
-HGA [15] 0.2586 0.1842 0.1085 0.0517 0.0365 0.0148 0.1433 0.0932 0.2909 0.1806 0.2908 0.2128 0.0877 0.0266 0.5231 0.4078 2.9111
-(b) (b)
-Model BLEU-1 BLEU-2 BLEU-3 METEOR ROUGE SPICE CIDEr S-BERT CapsMIX
-A E A E A E A E A E A E A E A E
-BlindQA [2] 0.2312 0.1847 0.0869 0.0550 0.0319 0.0198 0.1263 0.0930 0.2519 0.1806 0.2758 0.2447 0.0667 0.0265 0.4844 0.4153 2.7747
-UATT [51] 0.2659 0.1936 0.1180 0.0645 0.0548 0.0223 0.1423 0.0868 0.3025 0.2200 0.3449 0.2595 0.0940 0.0713 0.5447 0.4750 3.2601
-HME [6] 0.2640 0.1930 0.1238 0.0644 0.0556 0.0239 0.1457 0.0868 0.2998 0.2155 0.3115 0.1623 0.0944 0.0698 0.5341 0.4761 3.1207
-HGA [15] 0.2674 0.2293 0.1337 0.0823 0.0620 0.0308 0.1703 0.1123 0.3186 0.2363 0.3734 0.2558 0.1972 0.0477 0.5765 0.4692 3.5628
-(c) (c)
-Table 3: OEAG Results on our dataset. (a) UD split; (b) PS split; (c) UN split.
-4.2 Benchmarking on OEAG
+4.2 Benchmarking on Open-ended Answer Generation (OEAG)
 CapsMIX Performance metric.
-We measure the performance of open-ended generation in terms of BLEU-1,2,3 [33], METEOR [3], ROUGE [23], SPICE [1], CIDEr [44] and Sentence-BERT [36] scores by comparing with the ground truth answers as done by captioning and QA literature.
-However, we note that with such a wide range of metrics, it is difficult to get a comprehensive insight into models’ performances and compare them.
-To address that, we introduce a comprehensive metric, termed Caps-MIX (Captioning Metrics Integration eXpert), which integrates all the previously mentioned scores after normalizing them to their theoretical best values.
-This not only makes it more accessible to compare models based on just a single number but also combines the characteristics of individual metrics, each of which measures performance from a unique perspective.
-We note that the WUPS [27] score was originally developed for measuring the performance of single-word answers, & may not be an ideal metric for our dataset, which contains longer, detailed answers.
-Thus, we refrain from using it.
+We measure the performance of OEAG in terms of BLEU-1,2,3 [33], METEOR [3], ROUGE [23], SPICE [1], CIDEr [44] & Sentence-BERT [36] scores by comparing with the ground truth answers as done by captioning & QA literature.
+The wide range of metrics complicates model comparison, so we introduce Caps-MIX (Captioning Metrics Integration eXpert), which normalizes and integrates all scores into a single metric, simplifying comparisons and combining the unique strengths of individual metrics.
+Model CapsMIX
+BlindQA 2.7646
+UATT 3.3928
+HME 3.0975
+HGA 3.7872
+BlindGPT-2 6.6006
+VisionGPT-2 6.7582
+BLIP-2 1.8931
+Video-LLaMA 2.4464
+VideoChat2 3.9524
+GPT-4o 2.9851
+Table 3: OEAG results on our dataset (UD split).
 Baselines.
-Following prior work [47], we benchmarked the performance of EVQA [2]; UATT [51]; HME [6]; HGA [15] on open-ended answer generation task.
-Additionally, we also report the zero-shot performance of recent multimodal video understanding models BLIP-2 [18], Video-LLaMA [57], VideoChat2 [21].
+Following prior work [47], we benchmarked the performance of EVQA [2]; UATT [51]; HME [6]; HGA [15] on OEAG task.
+We also report the zeroshot performance of recent multimodal video understanding models BLIP-2 [18], Video-LLaMA [57], VideoChat2 [21], GPT-4o [31].
 We evaluated the baselines for generating answers and explanations.
-The performances of various models are presented in Table 3.
-Comparing the BLEU scores, we can see that the performances degrade as we move from BLEU-1 to BLEU-3, indicating that the models are struggling with generating longer answers.
-We confirmed this qualitatively.
+The performances of various models are summarized in Table 3.
+For full results, see Appendix.
+Overall, we observe that models struggle significantly with open-ended generation, including some recent VLMs and MLLMs.
+Despite GPT-4o’s strong performance on MCQA, it also falters on OEAG.
+These models likely perform better on MCQA by eliminating incorrect choices, but they fail to genuinely understand videos and perform causal reasoning for standalone answer generation.
+VideoChat2 seem to be doing exceptionally well, perhaps, because it was specifically trained on various video understanding tasks and datasets, including causal reasoning task.
 Comparing split-type wise, we found that as in the case of MCQA, UD was the easiest split.
-Overall, we conclude that models struggle quite a bit with open-ended generation.
-Inspired by the success of large language models (LLMs), we experimented with leveraging GPT-2 [35], a publicly available LLM, as our natural language answer generator.
+Inspired by the success of LLMs, we experimented with leveraging GPT-2 [35], a publicly available LLM, as our natural language answer generator.
 We found that a pre-trained version out of the box did not work well—it generated mostly random, unrelated words.
-However, upon simple training on the train set, it performed significantly better than the baselines we considered as shown in Table 3(a) (BlindGPT-2).
-However, it is unclear if these large language models have “seen" Tom and Jerry scripts during their pretraining stage.
+However, upon simple training on the train set, it performed significantly better than the baselines we considered as shown in Table 3 (BlindGPT-2).
+However, it is unclear if these LLMs have “seen" Tom and Jerry scripts during their pretraining stage.
 If so, then pretraining on relevant scripts followed by finetuning on our dataset could potentially be a reason for GPT-2’s good performance.
@@
 In the next step, we integrated visual information into LLM.
-Taking inspiration from [31], we learn a projector network to align the visual features with GPT-2 representations (VisualGPT-2 in Table 3(a)).
+Taking inspiration from [30], we learn a projector network to align the visual features with GPT-2 representations (VisualGPT-2 in Table 3.
 We observed an improvement in the performance.
@@
 + Ours 32.91 41.44 55.96 1.48
-Table 4: Our dataset improves performance on existing real-world dataset. For OEAG, CapsMIX score is reported.
+Table 4: Our dataset improves performance on existing real-world dataset. For OEAG, WUPS score is reported.
 4.3 Does our dataset help with real-world cases?
-To observe our dataset’s direct and immediate impact, we combine our training set with NextQA’s training set in this experiment.
-Then, we note the performance of models on NextQA test set.
-For comparison, we also note the performance of models when our dataset is not leveraged/incorporated.
-Despite being a synthetic/cartoon dataset, our dataset boosted the performance on a real-world dataset on both MCQA and OEAG (Table 4).
+To evaluate the direct impact of our dataset, we combine our training set with NextQA’s training set and measure model performance on the NextQA test set.
+For comparison, we also measure model performance without incorporating our dataset.
+Despite being a synthetic/cartoon dataset, and much smaller in size, our dataset boosted the performance on a real-world dataset on both MCQA and OEAG (Table 4).
 We observed improvement in identifying the correct causes by breaking the reliance on shortcuts and focus on causal effects; a more comprehensive analysis of the situation/interaction, rather than jumping to a conclusion.
-Qualitative results provided in Supplementary material.
+Qualitative results provided in Appendix.
 What is more, improvements were not limited to ‘causal-why’ questions but also extended across other question types.
@@
 Thus, we believe that it might be better to: 1) transfer reasoning skills acquired from challenging synthetic datasets like ours; 2) leverage our synthetic dataset to inform the model designing process, as it better reflects the challenges VideoQA models may face in the real world, such as longer causal chains and frequent scene changes in visual streams.
-Nonetheless, we do not recommend naively transferring or deploying whole models/weights trained on our dataset to real-world scenarios or applications.
+Nonetheless, we do not suggest naively deploying whole models/weights trained on our dataset to real-world scenarios or applications.
 5 Conclusion
-We introduced CausalChaos!, a challenging dataset for causal action question-answering tasks.
-Based on the timeless classic of the Tom and Jerry cartoon, it is richly annotated with critical thinking questions demanding a wide spectrum of reasoning from Video QA models.
-Questions are supported with multi-level answers and further explanations, all-encompassing video context.
-We further provide a novel CausalConfusion test set, which can challenge the causal relationship modeling capabilities of Video QA models.
-Through experimentation, we found that while existing models work well on causal action QA tasks, there is a lot of room to improve, especially, in terms of causal relationship modeling, and long and detailed open-ended answer generation.
-While LLMs offer a promising start, sophisticated integration of visual information with LLMs or joint modeling vision and language would be crucial.
-We hope that our challenging dataset will foster such developments.
-We will release our dataset and codes to help future efforts in this direction.
-Lastly, we showed that it brings improvement to real-world datasets.
-We believe our dataset, derived from cartoons, should inform the model designing process, as it better reflects the challenges Video-QA models may face in the real world, such as longer causal chains and frequent scene changes in visual streams.
-However, we do not suggest naively transferring or deploying whole models trained on our dataset to real-world scenarios.
+We introduced CausalChaos!, a challenging dataset for causal action question-answering tasks based on the classic Tom and Jerry cartoon, richly annotated with critical thinking questions requiring extensive reasoning from Video QA models.
+Questions come with multi-level answers and explanations covering the entire video context.
+We also provide the novel CausalConfusion test set to challenge causal relationship modeling in Video QA models.
+Our experiments show that while existing models perform well on causal action QA tasks, there is significant room for improvement in causal relationship modeling and generating detailed open-ended answers.
+LLMs show promise, but integrating visual information with LLMs or joint modeling of vision and language is crucial.
+We hope our dataset fosters such developments and will release it and the codes to support future efforts.
+Lastly, we demonstrated improvements in real-world datasets.
+Our dataset, derived from cartoons, should inform model design, reflecting real-world challenges like longer causal chains and frequent scene changes.
+However, we do not suggest deploying models trained solely on our dataset in real-world scenarios.
+Material Acknowledgement and Disclaimer.
+Tom and Jerry is a material of Turner Entertainment Company (Warner Bros.
+Entertainment Inc.).
+All rights reserved.
+We do not claim any ownership of or rights to the Tom and Jerry material.
+All other trademarks, service marks, trade names and any other material referenced in this document are the property of their respective owners.
 Acknowledgements.